OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of each should he use? Let p = pounds of peanuts and let c = pounds of cashews. Write a system of equations that could be used to solve the problem.

OpenStudy (anonymous):

3lbs of cashsews

OpenStudy (whpalmer4):

Cost of \(p\) pounds of peanuts is \(4p\) Cost of \(c\) pounds of cashews is \(6.5c\) We have two equations: \[p+c = 10\]We are making 10 pounds of the mixture \[4p+6.5c = 10*4.75\]Cost of 10 pounds of the mixture at the desired price of $4.75/lb. This is an unrealistic problem in that the poor grocer is expected to sell the mixture at cost! Margins are thin in the grocery business, but that's ridiculous!